OpenStudy (anonymous):

Describe the graph of the function f(x) = x3 − 18x2 + 107x − 210

OpenStudy (danjs):

what info do you usually need... this is a cubic equation, polynomial

OpenStudy (danjs):

remember the look of y=x^3 

OpenStudy (anonymous):

Describe the graph of the function f(x) = x3 − 18x2 + 107x − 210. Include the y-intercept, x-intercepts, and the shape of the grap

OpenStudy (anonymous):

Thats what I need

OpenStudy (danjs):

that is the general shape, the squared term will mess with that middle section a bit Graph the general form here y = A*x^3 + B*x^2+C*x+D Select 'add sliders' for ABCD, mess with those coefficient values and watch the graph

OpenStudy (danjs):

https://www.desmos.com/calculator

OpenStudy (danjs):

well the y intercept is when x=0, just solve that point

OpenStudy (anonymous):

I just dont know how to describe it, nor the shape

OpenStudy (danjs):

Maybe start by saying, the general shape is similar to that of y=x^3 did you make that graph with sliders on the other page?

OpenStudy (danjs):

notice if A is + vs A is -, the way the ends of the graph look

OpenStudy (anonymous):

Yeah made the graph on the page you sent me

OpenStudy (danjs):

You can show that a cubic has at least one real root (x-intercept), because if you take very large values for x, the cube term dominates, and the ends of the graph for x is large + and -, go off to huge values for f(x) in opposite directions... the range is then all real numbers, and it must cross the x axis somewhere at least once then, to go from -huge number to a + huge number, the sign must change

OpenStudy (danjs):

Go to that site, in the place where you put things in on the left.. put a*x^3+b*x^2+c*x+d hit to add sliders for all 4, a,b,c, and d play with the values and watch the effects

OpenStudy (anonymous):

Okay Thanks love

OpenStudy (danjs):

you can see how each term affects the graph, for yorus f(x) = x3 − 18x2 + 107x − 210 a=1 b=-18 c=107 d=-210 probably have to zoom way out and set the scale on the axis to see it good

OpenStudy (danjs):

just describe what it looks like and why

OpenStudy (danjs):

For huge x values, the function f(x) goes off to +infinity because the coefficient on the x^3 term 'A' is positive...

OpenStudy (danjs):

to get the x intercepts, (must be at least one real value), solve for f(x) =0 if it is a book problem, that probably factors nicely so you can get nice values for x

OpenStudy (danjs):

f(x) = x3 − 18x2 + 107x − 210 = ( )*( )* ( ) = 0 ( )*( )* ( ) = 0 you can tell from the graph with those values for B and C and D, it will have 3 x-intercepts, so it may factor into that form

OpenStudy (anonymous):

for those x - intercept wouldnt it be (5,0) (6,0) (7,0)

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

yes, that factors nice into those 3 quantities

OpenStudy (danjs):

for f(x) =0, x can be 5, or 6 or 7

OpenStudy (anonymous):

So for y it would be -210

OpenStudy (danjs):

yes, (0,y) , when x is zero

OpenStudy (danjs):

Maybe explain the end behavior for the graph when you let x be very large possitive and negative... the larger x becomes, the more the x^3 term will completely dominate the value of f(x)

OpenStudy (anonymous):

I just dont know how to put all that in an explanation does that make sense

OpenStudy (danjs):

for this one, as x becomes larger and larger, the values of f(x) also become larger and larger without bound, , same for -x, and f(x) going to huge - values

OpenStudy (danjs):

that explains the end behavior of the graph shooting upwards and downwards for ever to infinity

OpenStudy (danjs):

what class is this?

OpenStudy (anonymous):

obviously its a math class lol

OpenStudy (anonymous):

But for this last assignment I need to explain it but I swear I am horrible at explaining things

OpenStudy (anonymous):

So this is so hard for me,

OpenStudy (danjs):

i mena, algebra, pre-calc, calculus ?

OpenStudy (danjs):

If you have to write it out, just go throgh all the posts from me above and should be enough

OpenStudy (anonymous):

algebra 2